Outline and explain a put and call purchase strategy that will allow me to use $10,000 to control 100 shares of the common stock in my portfolio. I need specific examples of how I would do this as a strategy in my investment portfolio.  
Answer #1

No. of options contract in this case = 10,000/100

= 100 option contracts

  • In case of Long position (purchase), call option, the pay off is spot price - strike price.

Where, total profit is spot price - strike price - premium.

Suppose, a trader buys call option at a premium of $2 and enters in to a contract of 3 months call option at a exercise price of $20. Spot price at the time of maturity is $30.

So, profit for trader is 30-20-2= $8 per option. =8×100 =$800.

So, call option (purchase) is exercise only when spot price is greater than strike price or exercise price that is when the trader is expecting the rise in price at the time of maturity.

  • In case of put long position, the pay off is strike price minus spot price. Where, the total profit is strike price -spot price - premium.

Suppose, a trader buys put option at a premium of $3 and enters in to a contract of 3 months put option at exercise price of $28 and spot price at the time of maturity becomes $ 20.

So, profit for trader is 28 - 20 - 3 = $5 per option = 5×100 = $500

So, put option is exercise only when strike price is greater than spot price in case of Long position that is when trader is expecting fall in price at the time of maturity.